40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Apr 29 22:53:06 2018, gautam, Update, General, DARM actuation estimate darmEst.pdf
    Reply  Tue May 1 19:37:50 2018, gautam, Update, General, DARM actuation estimate darmEst.pdfdarmEst_time.pdf
       Reply  Mon May 7 16:23:06 2018, gautam, Update, General, DARM actuation estimate darmProj.pdf37.pngMCalign_20180507.png
          Reply  Mon May 7 20:01:14 2018, Rorpheus, Update, General, Use anti-dewhitening + show CARMA/DARMA 
             Reply  Thu May 10 14:13:22 2018, gautam, Update, General, More refinement of DARM control signal projection darmProj.pdfdarmOLTF.pdfDWcomparison.pdf
                Reply  Fri May 11 13:58:42 2018, rana, Update, General, More refinement of DARM control signal projection 
                   Reply  Fri May 11 19:02:52 2018, gautam, Update, General, More refinement of DARM control signal projection darmProj.pdfdarmOLTF.pdfDWcomparison.pdfDARMsuppression.pdfControlLoop.pdf
                      Reply  Sat May 12 10:02:03 2018, rana, Update, General, More refinement of DARM control signal projection 
                         Reply  Sun May 20 17:43:01 2018, rana, Update, Electronics, How to choose resistors 
                      Reply  Mon May 14 18:58:32 2018, Kevin, Update, PonderSqueeze, Squeezing with no SRM homodyne_heatmap.pdfdisplacement_noise.pdfnoise_budget.pdf
                         Reply  Tue Oct 2 23:57:16 2018, gautam, Update, PonderSqueeze, Squeezing scenarios PonderSqueeze_NB_LORIN.pdf
                   Reply  Tue May 15 21:56:57 2018, gautam, Update, General, Stack measurement setup decommissioned 
                      Reply  Thu May 17 09:14:38 2018, Steve, Update, General, Stack measurement setup decommissioned Load_Cell_Measurement_Set_Up.jpgETMX_stack_up_down.png
Message ID: 13799     Entry time: Sun Apr 29 22:53:06 2018     Reply to this: 13805
Author: gautam 
Type: Update 
Category: General 
Subject: DARM actuation estimate 

Motivation:

We'd like to know how much actuation is required on the ETMs to lock the DARM degree of freedom. The "disturbance" we are trying to cancel is the seismic driven length fluctuation of the arm cavity. In order to try and estimate what the actuation required will be, we can use data from POX/POY locks. I'd collected some data on Friday which I looked at today. Here are the results.

Method:

  • I collected the error and control signals for both arm cavities while they were locked to the PSL.
  • Knowing the POX/POY sensing response and the actuator transfer functions, we can back out the free running displacements of the two arm cavities.
    • I used numbers from the cal filters which may not be accurate (although POX sensing response which was recently measured).
    • But the spectra computed using this method seem reasonable, and the X and Y arm asds line up around 1 Hz (albeit on a log scale).
    • In this context, L_X is really a proxy for |f_X - f_{MC}| and similarly for L_Y so I think the algebra works out correctly.
    • I didn't include any of the violin mode/AA/AI filters in this calculation.
  • Having calculated the arm cavity displacements, I computed "DARM" as L_y- L_x and then plotted its asd.
  • For good measure, I also added the quadrature sum of 4 optics' displacement noise as per the 40m GWINC model - there seems to be a pretty large discrepancy, not sure why.

If this approach looks legit, I will compute the control signal that is required to stabilize this level of disturbance using the DARM control loop, and see what is the maximum permissible series resistance we can use in order to realize this stabilization. We can then compare various scenarios like different whitening schemes, with/without Barry puck etc, and look at coil driver noise levels for each of them. 

Attachment 1: darmEst.pdf  303 kB  Uploaded Sun Apr 29 23:57:11 2018  | Hide | Hide all
darmEst.pdf
ELOG V3.1.3-